Can someone please help me with this problem:
A company sells a product at $60 per unit that has unit variable costs of $40. The company's break-even sales volume is $120,000. How much profit will the company make if it sells 4,000 units?
a. $240,000
b. $40,000
c. $80,000

Answer by gonzo(654)

company breaks even at sales volume of $120,000
to make $120,000, it had to sell $120,000 / $60.00 revenue per unit sold.
number of units sold = $120,000 / $60.00 = 2000.
variable cost for 2000 units sold is 2000 * $40.00 = $80,000.
profit = revenue - cost
cost = variable cost plus fixed cost.
let p = profit
let r = revnue
let f = fixed cost
let v = variable cost
formula is:
p = r - (f + v)
remove parentheses to get:
p = r - f - v
since the company breaks even at a volume of $120,000 in revenue, this means that:
p = 0
r = $120,000
f = f
v = $80,000
and the formula becomes:
$0 = $120,000 - f - $80,000
which becomes:
$0 = $40,000 - f
add f to both sides of the equation to get:
f = $40,000
since you know the fixed cost, you can now determine the profit on sales of 4000 units.
formula is:
p = r - f - v
f = $40,000
v = $40.00 * 4000 = $160,000
r = $60.00 * 4000 = $240,000
formula becomes:
p = $240,000 - $40,000 - $160,000
which becomes:
p = $240,000 - $200,000
which becomes:
p = $40,000
company will make $40,000 profit if it sells 4000 units.